Grade 12th passPhysical Chemistry

Samples of F2 gas & Xe gas are mixed in a container of fixed volume. The partial pressure of F2 gas& Xe gas before they react are 8*10-5 kPa and 1.2*10-5kPa respectively. When all of the Xe gas has reacted forming a solid compound the partial pressure of the remaining F2 gas was 4.6*10-5 kPa. The system was maintained at a constant temperature during the above process. What is the formula of tje above compound formed?

Madeena
6 Years agoGrade 12th pass

1 Answer

Arun
6 Years ago
Answer is: XeF₄.
p₁ (F₂) = 8,0 atm.
p₁ (Xe) = 1,7 atm.
p₂ (F₂) = 4,6 atm.
p₂ (Xe) = 0 atm, all reacted.
p - partial pressure.
pressure of F₂ reacted = Δp (F₂) = 8.0 - 4.6 = 3.4 atm.
pressure of Xe reacted = Δp(Xe) = 1,7 - 0 = 1,7 atm.
Δp (F₂)  / Δp(Xe) = 3,4 atm / 1,7 atm = 2:1.
Reaction: 2F₂ + Xe → XeF₄ (xenon-tetrafluoride)